# How Much are Large Language Models Contaminated? A Comprehensive Survey and the LLMSanitize Library

### v3

@@ 3.1 Performance Analysis
We first review a line of research assessing model contamination through simple performance analysis. $M$ is applied to different evaluation datasets, and a sudden drop in performance suggests previous contamination.
The most straightforward performance analysis consists in plotting performance against time, which leverages the fact that evaluation datasets released after a LLM is pre-trained should not be contaminated. In the code domain, Roberts et al. (2023) investigate the performance of GPT-3.5 and GPT-4 on datasets released before and after the training cutoff for both models. The analysis underscores a statistically significant positive correlation for problems catalogued on GitHub prior to the GPT models’ training cut-off, suggesting that their inclusion in the training data enhanced their performance on related tasks. Concurrent to their work, Li and Flanigan (2023) also plot the performance of several LLMs including GPT-3 (Brown et al., 2020) and open-source LLMs on general natural language understanding tasks against time, specifically before and after each LLM’s training data cutoff date. They find that LLMs are much more likely to improve on a trivial majority baseline for datasets released before their cutoff date, suggesting some contamination. While evaluating on fresh data guarantees a contamination-free performance evaluation, it is hard to scale this method in time as it requires frequent update of evaluation sets. Besides, the work of Cao et al. (2024) shows that evaluating on data posterior to the LLM training cutoff time may not always show a decrease in performance: in the code domain, the opposite may happen, complicating the usage of training cutoffs as a contamination detection method.
@@
Jiang et al. (2024b) delve deep into the impact of evaluation data leakage within the pre-training set by running a series of controlled pre-training experiments with GPT-2 (Radford et al., 2019). They pre-train from scratch GPT-2 in three different ways: with the original pre-training set, with additional input texts from four evaluation datasets (SST-2, MMLU (Hendrycks et al., 2020), CNN/DM (Hermann et al., 2015) and SQuAD-v1 (Rajpurkar et al., 2016)), and with input texts and labels from these same evaluation datasets. Contaminated pre-training clearly boosts performance on all downstream tasks. Findings also reveal a U-shaped relationship between the number of times that contaminated data points are included in the pre-training, and model performance. Authors also show that the n-gram contamination detection method from Llama-2 Touvron et al. (2023) is not effective at detecting contaminated data points, calling for more robust methods.
@@ 3.2 Model Memorization
An intuitive approach to detect contamination in LLMs is to frame it as a memorization problem. Researchers focus on which training data points have been memorized verbatim by the model.
Historically, this problem has been addressed in the field of privacy-preserving machine learning: adversarial attacks were crafted as membership inference attacks to access training data points memorized by a black-box language model, potentially compromising sensitive private information. We refer the reader to relevant literature in this research direction (Carlini et al., 2021, 2022).
Elangovan et al. (2021) highlight that some LLMs may benefit from an overlap between training and test sets, and merely memorize data points in this intersection, inflating their test set performance. This overlap is due to the common practice of simple random data shuffling in NLP tasks. By employing a bag-of-words approach to assess text similarity, the study provides a foundational method for identifying and mitigating leakage, though it acknowledges the complexity of semantic similarity measurement and the potential for more sophisticated methodologies to refine overlap detection. Magar and Schwartz (2022) find evidence of contamination across several benchmarks, impacting the performance of models like GPT-3. They propose to quantify the effect of contamination by training on a mix of general and task-specific data, then comparing performance on seen versus unseen instances to measure memorization and exploitation. This methodology highlights the nuanced relationship between memorization and exploitation. Gemma Team Google DeepMind (2024) emphasize the potential vulnerabilities of aligned models to adversarial attacks inducing recitation of memorized training data. Through comparable methods to Anil et al. (2023), they analyze discoverable memorization across 10k documents from the Gemma pre-trained models’ corpora. The analysis distinguishes between exact memorization — where the model’s output precisely matches the source — and approximates memorization, gauged by a 10% edit distance threshold. Gemma models exhibit memorization rates comparable to those of the PaLM and PaLM-2 models.
@@
Cloze tasks constitute a special type of prompting where parts of the input is masked. Chang et al. (2023) design a cloze task in which the LLM is prompted with a book’s passage, from which the character name has been replaced by a [MASK] token. Correctly predicting the character’s name out of several passages per book assesses whether the LLM has seen this book during pre-training. The name cloze accuracy is very high for GPT-4 on several popular books such as Alice’s Adventures in Wonderland, and shows a striking difference with BERT’s performance on the same task. Deng et al. (2023) present the TS-Guessing method: one of the options in question-answering benchmarks test sets is hidden, and the LLM is asked to predict it. Exact match rates above 50% for OpenAI models suggest contamination on the MMLU dataset (Hendrycks et al., 2020). The work of Ranaldi et al. (2024) also uses a cloze task: they prompt the LLM to reconstruct columns names in SQL dumps, and compare the performance on an older and more recent dataset to assess contamination. GPT-3.5 accuracy falls from 33.42% to 13.21%, indicating contamination on the older Spider dataset.
@@ 3.3 Model Confidence
Going beyond simple string recitation of memorized data points, researchers may analyze the LLM’s confidence to detect contamination: a (too) high confidence may signify prior exposure to the data. Confidence analysis yields several powerful contamination detection techniques, which we highlight require white-box access to the LLM, as researchers need to obtain the output probability distribution for each token during inference.
@@
Next-token likelihood and perplexity (by extension to a whole dataset) are widely used in contamination detection. Given a test sample $X$, the Min-KProb technique (Shi et al., 2024) consists in running the LLM through all tokens of $X$, and keeping track of the $k$tokens with the smallest predicted probability. Then, the average between such bottom probabilities is computed and $X$ is deemed contaminated if this average is too high. Li (2023a) also work on token-level probabilities and compare perplexity on benchmark samples against memorized and clean baselines. The study finds significant memorization in recent models on popular reading comprehension and summarization benchmarks, while multiple-choice benchmarks show less evidence of contamination. This method provides a tool for the community to conduct rigorous contamination analysis, enabling more accurate and reliable model evaluation. Dong et al. (2024) propose two novel likelihood-based contamination detection methodologies: CDD (Contamination Detection via output Distribution) and TED (Trustworthy Evaluation via output Distribution). CDD detects data contamination by observing the peakedness in the LLM’s output distribution in a black-box manner. It represents a significant improvement over existing approaches, offering average relative improvements of 21.8%-30.2% in terms of Accuracy, F1 Score, and AUC metrics. TED corrects the LLM’s output distribution to mitigate the impact of data contamination on evaluation metrics, significantly reducing performance improvements attributed to data contamination across various scenarios and contamination degrees. The work of Xu et al. (2024b) proposes to use perplexity and accuracy on next n-gram prediction. They compute the relative loss in performance on these metrics on paraphrased versions of the dataset created by ChatGPT compared to the original dataset. The difference of relative loss between training and test set is later used to assess contamination or benchmark leakage. Experiments on GSM8K (Cobbe et al., 2021) hint at contamination for QWen (Bai et al., 2023), Aquila and InternLM (Cai et al., 2024) series of LLMs.
@@
Oren et al. (2023) take an original approach to analyze the LLM’s likelihood on an evaluation dataset $D_{E}$. They run inference on $D_{E}$ and on shuffled versions of $D_{E}$. LLM log probabilities being statistically different on the non-shuffled dataset compared to the shuffled versions signifies contamination. The method relies on the assumption that if they are included in the pre-training set, evaluation datasets tend to be present with the same default ordering.
3.4 Model-based Closed-Data Contamination Detection
Relying on an external, high-performing LLM like GPT-4 can be useful in closed-data contamination detection, similarly to the open-data setup.
Golchin and Surdeanu (2023b) make use of GPT-4’s abilities in guided prompting, an enhanced prompting process where the completion prompt includes extra information such as the dataset name. Contamination is then assessed based on the averaged difference in performance between standard and guided prompting, or if GPT-4 with in-context learning finds an exact match or two near-exact matches in the guided completions. This latter method with GPT-4 is highly accurate in identifying contamination (92%–100% accuracy). Moreover, the investigation highlights the prevalence of contamination in datasets such as AG News (Zhang et al., 2015), WNLI (Levesque et al., 2012), and XSum (Narayan et al., 2018), emphasizing the critical need to address data integrity in LLM applications. The same authors follow up with the Data Contamination Quiz (DCQ) evaluation framework for contamination detection in black-box LLMs (Golchin and Surdeanu, 2023a). LLMs are prompted with five completion options, where one option is the ground-truth text from the original dataset, three other options are paraphrases from GPT-4, and the last choice is None. The assumption is that if the LLM picks the exact answer, it is doing so out of memorization, and the authors show that this DCQ framework finds contamination cases than the guided prompting method.

### v4

@@ 3 Closed-Data Contamination Detection
There are two broad ways to assess closed-data contamination: examining model behavior aggregated on the whole dataset, and focusing on instance-level outputs. For the former, the main method, which we explore in Section 3.1, consists in tracking model performance on several datasets, sampled from different distributions, notably through time. For the latter, we review methods ranging from the strongest assessment of contamination to the weakest: membership-inference attacks, which assess the presence of an entire data point in the pre-training set (Section 3.2) ; then memorization of spans of text by the model (Section 3.3) ; and lastly methods analyzing model confidence on specific tokens (Section 3.4).
@@ 3.1 Performance Analysis
We first review a line of research assessing model contamination through simple performance analysis. $M$ is applied to different evaluation datasets, and patterns in evaluation scores may indicate contamination.
3.1.1 Train-Test Performance Gap
On a dataset where the training set and the test set are sampled from the same distribution, a machine learning system is expected to perform on the test set slightly worse, or at best similar to, the training set. Zhao et al. (2024) leverage this fact to distinguish between standard fine-tuning and contamination, focusing on domain-specific setups such as financial data. They introduce a Performance Consistency Ratio (PCR), equal to the LLM performance on the test set divided by its consistency, where the latter measures the LLM’s ability to produce the same output across several reasoning paths. A PCR on the training set significantly greater than on the test set is typical of fine-tuning ; the opposite case points out at contamination. Findings include potential contamination of Baichuan-13B (Yang et al., 2023a) on FinEval (Zhang et al., 2023) and of the FinMA model (Xie et al., 2023) on the FinQA dataset (Chen et al., 2021b).
3.1.2 Controlled Training Setup
@@
Jiang et al. (2024b) delve deep into the impact of evaluation data leakage within the pre-training set by running a series of controlled pre-training experiments with GPT-2 (Radford et al., 2019). They pre-train from scratch GPT-2 in three different ways: with the original pre-training set, with additional input texts from four evaluation datasets (SST-2, MMLU (Hendrycks et al., 2020), CNN/DM (Hermann et al., 2015) and SQuAD-v1 (Rajpurkar et al., 2016)), and with input texts and labels from these same evaluation datasets. Contaminated pre-training clearly boosts performance on all downstream tasks. This was also shown by the work of Palavalli et al. (2024) in their Verbatim setup when pre-training GPT-2. Findings also reveal a U-shaped relationship between the number of times that contaminated data points are included in the pre-training, and model performance. Authors also show that the n-gram contamination detection method from Llama-2 Touvron et al. (2023b) is not effective at detecting contaminated data points, calling for more robust methods.
3.1.3 Performance Change Through Time
The most straightforward performance analysis consists in plotting performance against time, which leverages the fact that evaluation datasets released after a LLM is pre-trained are new and should not be contaminated.
Aiyappa et al. (2023) are among the first to question the non contamination of ChatGPT training data by popular evaluation benchmarks. Their run of the January 30th, 2023 version of ChatGPT on the SemEval 2016 dataset shows a 12.5 macro-F1 score improvement compared to a prior experiment with the v1 of ChatGPT, released two months prior in November 2022. In the code domain, Roberts et al. (2023) investigate the performance of GPT-3.5 and GPT-4 on datasets released before and after the training cutoff for both models. The analysis underscores a statistically significant positive correlation for problems catalogued on GitHub prior to the GPT models’ training cut-off, suggesting that their inclusion in the training data enhanced their performance on related tasks. Concurrent to their work, Li & Flanigan (2023) also plot the performance of several LLMs including GPT-3 (Brown et al., 2020) and open-source LLMs on general natural language understanding tasks against time, specifically before and after each LLM’s training data cutoff date. They find that LLMs are much more likely to improve on a trivial majority baseline for datasets released before their cutoff date, suggesting some contamination. While evaluating on fresh data guarantees a contamination-free performance evaluation, it is hard to scale this method in time as it requires frequent update of evaluation sets. Besides, the work of Cao et al. (2024) shows that evaluating on data posterior to the LLM training cutoff time may not always show a decrease in performance: in the code domain, the opposite may happen, complicating the usage of training cutoffs as a contamination detection method.
3.2 Membership-Inference Attacks
Extracting knowledge from a language model has traditionally been addressed in the field of privacy-preserving machine learning. Adversarial attacks were crafted as membership inference attacks (MIA) to access training data points memorized by a black-box language model, potentially compromising sensitive private information (Carlini et al., 2021; 2022). If there is access to data points which are known not to be present in the training set, membership inference attach is referred to as reference-based MIA ; other cases belong to the setup of reference-free MIA.
3.2.1 Reference-based Membership-Inference Attacks
Xie et al. (2024) derive a MIA named RECALL based on the conditional language capacity of LLMs. RECALL conditions with non-member prefixes (for instance, sampled after pre-training cutoff date): log-likelihood drops more on average for member data points than for non-members, yielding a simple yet powerful MIA. The method reaches state-of-the-art ROC-AUC on the WikiMIA dataset (Shi et al., 2024). Meeus et al. (2024a) revisit MIA in the era of LLMs and define document-level MIA, which is the task of detecting whether an entire document (book, scientific paper, etc) has been trained on. A reference set is built by sampling documents known to be part of the pre-training set (e.g. Common Crawl), and data posterior to the pre-training cutoff date, which is guaranteed to be uncontaminated. Then, the authors derive features from predicted token probabilities by the LLM, which are normalized in several fashions. A meta-classifier predicts membership inference, reaching ROC-AUC as high as 0.86.
In the last stage of LLM post-training, PPO (Schulman et al., 2017) or DPO (Rafailov et al., 2023) is usually employed to train LLM on preference data. Feng et al. (2024) show that LLMs are vulnerable to MIA on the preference data, especially when using DPO. LLMs themselves can be instruction-tuned to detect if a data point is a member of their pre-training data. MIA-Tuner (Fu et al., 2024a) fine-tunes a soft prompt (Lester et al., 2021) on several target open-source LLMs to classify membership inference, and outperforms all other methods on WikiMIA. Even though at inference time, non-members are not needed, the methodology should still be assessed as a reference-based MIA due to the usage of non-members during the LLM fine-tuning.
3.2.2 Reference-free Membership-Inference Attacks
Fu et al. (2023) argue that both reference-based and reference-free MIA are unsatisfying with LLMs in practical scenarios. The latter rely on overfitting to assess membership inference. In the reference-free setup, they propose to prompt the target LLM to generate pseudo reference data. Semantically similar and semantically different paraphrases of target records are generated, and a membership signal is designed by measuring and comparing probabilistic variations between the target and reference models. This method reaches a ROC-AUC of 0.95 with Llama on the language modelling dataset WikiText-103. Chang et al. (2024) analyze the dynamics of next-token prediction loss for open-source Pythia and GPT-Neo LLM series. They observe that for members, the loss tends to be smoother and to decrease as context increases, whereas non-members loss curves show a flat slope as well as quite frequent outlier values. From this observation the authors derive a state-of-the-art MIA technique called CAMIA. EM-MIA (Kim et al., 2024) introduces a MIA method iteratively refining the membership score used to assess MIA and the prefix score through an expectation-maximization algorithm. The prefix score measures how discriminative the data point is in classifying members and non-members when used as a prefix. EM-MIA reaches state-of-the-art ROC-AUC scores on WikiMIA, however on a custom benchmark where members and non-members distributions are almost identical, performance drops back to random range, as for all other reported MIA methods.
3.2.3 Limitations of Membership-Inference Attacks
Membership-Inference attacks are hard in practice with neural networks. Duan et al. (2024) evaluate a series of MIAs on the Pythia LLMs (from 70M to 12B parameters), pre-trained on the Pile. Across types of attacks and domains, performance hardly surpasses random guessing (ROC-AUC inferior to 0.6). This failure is attributed to the common practice in LLM development of pre-training only for a single epoch and on a too large dataset, and the fact that the border between members and non-members is not clear. By splitting both groups based on time, where non-members are drawn from datasets posterior to the pre-training cutoff date, MIA accuracy drastically improves. This finding motivates the monitoring of LLMs performance through time described above. Meeus et al. (2024b) echo these findings from Duan et al. (2024): they train a trivial bag-of-words model on MIA datasets constructed by splitting on time and reach a high ROC-AUC. Authors argue that proper MIA evaluation should be done on a randomized train-test split. Concurrent work from Maini et al. (2025) also question the validity of membership inference attacks on a time-split of Wikipedia. An experiment replacing member sentences by sentences from the validation set of The Pile (Gao et al., 2020) using Pythia models (Biderman et al., 2023) shows membership inference tests still reaching an ROC-AUC of 0.7 ; proving that they merely capture a temporal distribution shift rather than actual membership to the pre-training set. Despite this lack of success of MIA on pre-training sets of LLMS, Li et al. (2024) show that similarity between the target sample and the generated output of a LLM-powered RAG system yields a powerful MIA on the RAG database. On Natural Questions (Kwiatkowski et al., 2019) and TriviaQA (Joshi et al., 2017), the method reaches more than 0.85 in both ROC-AUC and PR-AUC.
@@ 3.3 Model Memorization
An intuitive approach to detect contamination in LLMs is to frame it as a memorization problem. While MIAs assess the presence of an entire data point into the training set, here memorization is defined in a larger sense and focuses on detecting memorized sequences of tokens which may not be full data points (documents), but rather subsets of them.
3.3.1 Quantifying Memorization
Elangovan et al. (2021) highlight that some LLMs may benefit from an overlap between training and test sets, and merely memorize data points in this intersection, inflating their test set performance. This overlap is due to the common practice of simple random data shuffling in NLP tasks. By employing a bag-of-words approach to assess text similarity, the study provides a foundational method for identifying and mitigating leakage, though it acknowledges the complexity of semantic similarity measurement and the potential for more sophisticated methodologies to refine overlap detection. Magar & Schwartz (2022) find evidence of contamination across several benchmarks, impacting the performance of models like GPT-3. They propose to quantify the effect of contamination by training on a mix of general and task-specific data, then comparing performance on seen versus unseen instances to measure memorization and exploitation. This methodology highlights the nuanced relationship between memorization and exploitation.
The team from Google developing the Gemma model series (Gemma Team Google DeepMind, 2024) emphasizes the potential vulnerabilities of aligned models to adversarial attacks inducing recitation of memorized training data. Through comparable methods to Anil et al. (2023), they analyze discoverable memorization across 10k documents from the Gemma pre-trained models’ corpora. The analysis distinguishes between exact memorization — where the model’s output precisely matches the source — and approximate memorization, gauged by a 10% edit distance threshold. Gemma models exhibit memorization rates comparable to those of the PaLM and PaLM-2 models.
3.3.2 Prompt Engineering
@@
Golchin & Surdeanu (2023b) make use of GPT-4’s abilities in guided prompting, an enhanced prompting process where the completion prompt includes extra information such as the dataset name. Contamination is then assessed based on the averaged difference in performance between standard and guided prompting, or if GPT-4 with in-context learning finds an exact match or two near-exact matches in the guided completions. This latter method with GPT-4 is highly accurate in identifying contamination (92%–100% accuracy). Moreover, the investigation highlights the prevalence of contamination in datasets such as AG News (Zhang et al., 2015), WNLI (Levesque et al., 2012), and XSum (Narayan et al., 2018), emphasizing the critical need to address data integrity in LLM applications. The same authors follow up with the Data Contamination Quiz (DCQ) evaluation framework for contamination detection in black-box LLMs (Golchin & Surdeanu, 2023a). LLMs are prompted with five completion options, where one option is the ground-truth text from the original dataset, three other options are paraphrases from GPT-4, and the last choice is None. The assumption is that if the LLM picks the exact answer, it is doing so out of memorization, and the authors show that this DCQ framework finds more contamination cases than the guided prompting method. Samuel et al. (2024) also propose a quiz framework to detect contamination. Their Local Order Quiz method consists in prompting the LLM with a data point and asking it to select the subsequent data point in the original data set, among four possibilities shown in the prompt (the correct next data point, and three other random ones). A higher accuracy that the baseline 25% chance indicates that the LLM has memorized the data set. In practice, the IMDB (Maas et al., 2011) dataset shows high signs of contamination (subsequent data point selection accuracy over 80%) for all three LLMs under study: GPT-4 (Achiam et al., 2023), Claude-3  [^1] https://www.anthropic.com/news/claude-3-family and Llama-3 (Dubey et al., 2024).
3.3.3 Cloze Tasks
Cloze tasks constitute a special type of prompting where parts of the input is masked. Chang et al. (2023) design a cloze task in which the LLM is prompted with a book’s passage, from which the character name has been replaced by a [MASK] token. Correctly predicting the character’s name out of several passages per book assesses whether the LLM has seen this book during pre-training. The name cloze accuracy is very high for GPT-4 (Achiam et al., 2023) on several popular books such as Alice’s Adventures in Wonderland, and shows a striking difference with BERT’s (Devlin et al., 2019) performance on the same task. Deng et al. (2023) present the TS-Guessing method: one of the options in question-answering benchmarks test sets is hidden, and the LLM is asked to predict it. Exact match rates above 50% for OpenAI models suggest contamination on the MMLU dataset (Hendrycks et al., 2020). The work of Ranaldi et al. (2024) also uses a cloze task: they prompt the LLM to reconstruct columns names in SQL dumps, and compare the performance on an older and more recent dataset to assess contamination. GPT-3.5 accuracy falls from 33.42% to 13.21%, indicating contamination on the older Spider dataset (Yu et al., 2018).
@@ 3.4 Model Confidence
Going beyond simple recitation of memorized token sequences, researchers may now analyze the LLM’s confidence to detect contamination: a (too) high confidence on some tokens may signify prior exposure to the data. Confidence analysis yields several powerful contamination detection techniques, which we highlight require gray-box to white-box access to the LLM, as researchers need to obtain the output probability distribution for each token during inference.
3.4.1 Likelihood
Next-token likelihood and perplexity (by extension to a whole dataset) are widely used in contamination detection.
Given a test tokens sequence $X$, the Min-KProb technique (Shi et al., 2024) consists in running the LLM through all tokens of $X$, and keeping track of the $K$tokens with the smallest predicted probability. Then, the average between such bottom probabilities is computed and $X$ is deemed contaminated if this average is too high. In mathematical terms, noting $(x_{1},\ldots,x_{n})$ the sequence of evaluation tokens under study, Min-KProb is computed as :
[…]
A significant drawback is the choice of value for the hyper-parameter $K$, which authors recommend to set to 20 as default. On WikiMIA, Min-KProb reaches an average ROC-AUC of 0.72, and is the highest for every LLM tested. Notably, it outperforms other methods such as perplexity or the zlib compression entropy membership-inference attack. Due to its simplicity and effectiveness, Min-KProb is gaining traction as one of the most popular contamination detection methods. Zhu et al. (2024) use Min-KProb as a first contamination detection step to identify leaked test samples before re-writing them, in order to build cleaner versions of popular evaluation datasets GSM8K (Cobbe et al., 2021) and MMLU (Hendrycks et al., 2020). Min-KProb performance is very high ; yet the method has a major limitation: it does not account for the token frequency distribution. LLMs put a higher probability on more frequent tokens, thus a non-contaminated sequence made of only frequent tokens might get assigned a score higher than a sequence from the training set. To fix this, DC-PDD (Zhang et al., 2024b) calibrates token probabilities and computes the cross-entropy between the token probability distribution and the token frequency distribution. The latter is estimated through a large-scale publicly available reference corpus. The tokens sequence is scored by averaging these token-level cross-entropy scores over the first occurence of each token. Contamination is assessed when this final score is above a pre-defined threshold. Min-K%++ (Zhang et al., 2024a) proposes another extension of the idea of Min-KProb: rather than simply getting next-token predicted probabilities, the authors propose a score which subtracts the expected log-probability and divides the score by the variance; in other words, normalizing the initial log-probability. The sequence scoring mechanism is then identical to the one of Min-KProb. Formally, noting the expected probability over the next token $\mu_{x<i}=\mathbb{E}_{z\sim p(.|x_{<i})}[\text{log}(p(z|x_{<i}))]$ and its corresponding standard deviation $\sigma_{x<i}=\sqrt{\mathbb{E}_{z\sim p(.|x_{<i})}[(\text{log}(p(z|x_{<i}))-\mu_{x<i})^{2}]}$:
[…]
The motivation of the score built by Min-K%++ is to assess whether an input form a mode, after observing that because of maximum likelihood training, training samples often become local maxima in the modeled distribution along each input dimension (here in the context of LLM, dimensions are tokens). Min-K%++ reaches state-of-the-art on WikiMIA, outperforming Min-KProb by up to 10 points.
@@ 3.4.1 Likelihood
Li (2023a) also work on token-level probabilities and compare perplexity on benchmark samples against memorized and clean baselines. The study finds significant memorization in recent models on popular reading comprehension and summarization benchmarks, while multiple-choice benchmarks show less evidence of contamination. This method provides a tool for the community to conduct rigorous contamination analysis, enabling more accurate and reliable model evaluation. Dong et al. (2024) propose two novel likelihood-based contamination detection methodologies: CDD (Contamination Detection via output Distribution) and TED (Trustworthy Evaluation via output Distribution). CDD detects data contamination by observing the peakedness in the LLM’s output distribution in a black-box manner. It represents a significant improvement over existing approaches, offering average relative improvements of 21.8%-30.2% in terms of Accuracy, F1 Score, and AUC metrics. TED corrects the LLM’s output distribution to mitigate the impact of data contamination on evaluation metrics, significantly reducing performance improvements attributed to data contamination across various scenarios and contamination degrees. The work of Xu et al. (2024b) proposes to use perplexity and accuracy on next n-gram prediction. They compute the relative loss in performance on these metrics on paraphrased versions of the dataset created by ChatGPT compared to the original dataset. The difference of relative loss between training and test set is later used to assess contamination or benchmark leakage. Experiments on GSM8K (Cobbe et al., 2021) hint at contamination for Qwen (Bai et al., 2023), Aquila and InternLM (Cai et al., 2024) series of LLMs. Zhang & Wu (2024) leverage the concept of surprising tokens: these are tokens where the LLM is very confident yet wrong in its prediction of the next token. The authors define a score named SURP which is the average predicted log-probability on the ground-truth token for tokens of the sequence where the model shows low entropy (and therefore, high confidence) and where this log-probability is relatively smaller. A higher SURP score is thresholded to assess contamination. The method slightly outperforms Min-KProb on WikiMIA.
@@
Oren et al. (2023) take an original approach to analyze the LLM’s likelihood on an evaluation dataset $D_{E}$. They run inference on $D_{E}$ and on shuffled versions of $D_{E}$. LLM log probabilities being statistically different on the non-shuffled dataset compared to the shuffled versions signifies contamination. The method relies on the assumption that if they are included in the pre-training set, evaluation datasets tend to be present with the same default ordering. We point out that detecting the LLM’s memory of a dataset canonical order is an idea also explored by Samuel et al. (2024) for a different contamination detection method.
3.4.2 Probing Hidden Layers